OpenStudy (anonymous):

A street light is at the top of a 12 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 7 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 40 ft from the pole? Speed of tip of her shadow = ___ ft/sec

OpenStudy (anonymous):

What level math is this?

OpenStudy (anonymous):

Calculus.

OpenStudy (anonymous):

Derivatives and related rates.

OpenStudy (anonymous):

Oh sorry I'm not in calculus..

OpenStudy (anonymous):

Oh I just did those....this should help, its the same question with one different numebr :) http://openstudy.com/study#/updates/4f949b60e4b000ae9ecad05f

OpenStudy (anonymous):

It's okay.

OpenStudy (anonymous):

number*

OpenStudy (zale101):

i think u need to use v=dX/dt to identify the speed of the tip of her shadow

OpenStudy (zale101):

that's the the instantaneous velocity

OpenStudy (anonymous):



OpenStudy (anonymous):

\[{18\over{d+s}}={6\over s}\] which should give you \(18s=6d+6s\)

OpenStudy (anonymous):

now derive that and then i believe (not so sure, but-) that \(dd\over dt\) is supposed to be the 7 ft/sec since she's walking away -distance from the pole is increasing at the rate of- at a speed of 7 ft/sec

OpenStudy (anonymous):

\[18s=6d+6s\]\[18{ds\over dt}=6{dd\over dt}+6{ds\over dt}\]\[{ds\over dt}(18-6)=6(7)\]^ i pulled out the \(ds\over dt\) from the 18 and 6 after i subtracted the 6\(ds\over dt\) and then plugged in 7 ft/sec for the \(dd\over dt\)\[{ds\over dt}={6(7)\over(18-6)}\]\[{ds\over dt}={42\over12}\]\[{ds\over dt}={7\over2}ft/sec\]

OpenStudy (anonymous):

I'm really not sure if this is supposed to be the final answer @Jwoodward48 - sorry. But, i know the work I've done so far is correct...I hope this helps, and if it's possible, check the answer out in like the back of your book or something, or ask a friend what they got Sorry i couldnt help out too much

OpenStudy (anonymous):

It isn't correct, but thanks for trying.

OpenStudy (anonymous):

what's the actual answer?

OpenStudy (anonymous):

I'm still trying to figure it out.

OpenStudy (anonymous):

then how do you know it's wrong o.O

OpenStudy (anonymous):

the only other thing i can think of is the fact that you probably have to add her speed to the speed of her shadow...so that would be \({7\over2}+7\)\[={21\over2}ft/sec\]

OpenStudy (zale101):

i think she got the answer correct

OpenStudy (anonymous):

@Jwoodward48 the answer is \({21\over2}ft/sec\)

OpenStudy (anonymous):

I get the feeling you're not satisfied @Jwoodward48

OpenStudy (zale101):

he has to be

OpenStudy (zale101):

are u alright there bro?

OpenStudy (anonymous):

incorrect...

OpenStudy (anonymous):

It was 14!

OpenStudy (anonymous):

how do you know it's wrong...i'm confused and how is it 14?

OpenStudy (anonymous):

d=s

OpenStudy (anonymous):

Because of related rates.

OpenStudy (anonymous):

....i give up...for a second there i actually thought i was smart again..oh well

OpenStudy (anonymous):

12(l)=6(x+l) (you are smart, you used math to get an answer)

OpenStudy (anonymous):

12l-6l=6x

OpenStudy (anonymous):

6x=6l

OpenStudy (anonymous):

x=l

OpenStudy (anonymous):

dx/dt=dl/dt

OpenStudy (anonymous):

dx/dt=7

OpenStudy (anonymous):

I want dl/dt+dx/dt, which is 2 * dx/dt

OpenStudy (anonymous):

=14!

OpenStudy (zale101):

next time, you already know what u should be doing we learn from our mistakes D:

OpenStudy (anonymous):

Yep!

OpenStudy (zale101):

i took calculus last year, so i really don't have a complete understanding of it